If possible please give the formula or steps required for me to solve these questions, as I would like to go through the steps myself rather than just recieving answers. Thanks.

PART 1
Mass of Weighing Paper: 0.3966g
Mass of Weighing Paper & Oxalic Acid: 0.4705g
Initial Buret Reading: 0.00mL
Final Buret Reading: 12.80mL

1.) Moles of Oxalic Acid used?
2.) Protons (H+) avalivable for reaction with hydroxide ion (OH-)?
3.) Moles of hydroxide ion (OH-) which reacted with protons (H+)?
4.) Precise Molarity of NaOH solution?

PART 2
Proposed mass of unknown acid: 0.189g
Number of reactive protons per molecule: 2
Mass of weighing paper: 0.3958g
Mass of weighing paper and unknown acid: 0.4751g
Initial Buret Reading: 0.00mL
Final Buret Reading: 10.80mL

1.) Moles of NaOH which reacted with unknown acid?
2.) Moles of unknown acid which reacted with hydroxide ion (OH-)?
3.) Molecular weight of your unknown acid?
Answer #1

ANSWER - Part I:

a) Moles of oxalic acid used?

  • To calculate the moles of oxalic acid we use this formula:

muss moles = Molecular Weight

  • The molecular weight of oxalic acid (C2H2O4) is 90.03 g/mol
  • The mass of oxalic acid used is

mass C H204 = (Weighing paper & oralic acid) - (Weighing paper)

mass C2H204 = (0.47059) - (0.3966 g) = 0.0739 g

  • Then, the moles of oxalic acid used are

0.07399 moles = 90.03 g/mol = 8.21x10-4 mol

b) Protons (H+) available for reaction with hydroxyde ion (OH-)?

  • The oxalic acid (C2H2O4) ihas the following structure:.

ОН Но-

  • It is a dicarboxilic acid, this means that each molecule of oxalic acid has two protons available to react with OH-. Then, 1 mol of oxalic acid has 2 moles of protons:

1 mol CH204 2 mol H+

8.21x10-4 mol C H204 = 1.64x10-3 mol H+

c) Moles of hydroxide ion (OH-) which reacted with protons (H+)?

  • A neutralzation reaction has the following equation:

H+ + OH + H2O

  • This means that 1 mol of protons reacts with 1 mol of hydroxyde ion. Then, the moles of OH- used are:

1 mol H+ = 1 mol OH

1.64.r10-3 mol H+ +1.64x10-3 mol OH

d) Precise molarity of NaOH solution?

  • The molarity of solution is defined as

moles solute Molarity = Volume solution(L)

  • The moles of NaOH are

1 mol OH → 1 mol NaOH

1.64.r10-3 mol OH →1.64x10-3 mol NaOH

  • The volume of NaOH solution is

VNaOH = Final reading - Initial reading

V_{NaOH}=12.80\, mL-0.00\, mL=12.80\, mL

V_{NaOH}=\mathbf{12.80x10^{-3}\, L}

  • Then, the molarity NaOH solution is

Molarity=\frac{1.64x10^{-3}\, mol}{12.80x10^{-3}\, L}=\mathbf{0.128\, M}

---------------------------------------------------------------------------------------------------------------------------------------------------------------------

ANSWER - Part II:

a) Moles of hydroxide ion (OH-) which reacted with unknown acid?

  • We know the volume of 0.128 M NaOH solution used for reaction with unknown acid:

VNaOH = Final reading - Initial reading

V_{NaOH}=10.80\, mL-0.00\, mL=10.80\, mL

V_{NaOH}=\mathbf{10.80x10^{-3}\, L}

  • Now, we can calculate the moles of NaOH used

moles\, NaOH=M_{NaOH}\times V_{NaOH}

moles\, NaOH=0.128\, \frac{mol}{L} \times 10.80x10^{-3}L=\mathbf{1.38x10^{-3}\, mol}

b) Moles of unknown acid which reacted with hydroxide ion (OH-)?

  • Each mol of OH- (it comes from NaOH) reacts with one mol of H+ (it comes from unknown acid)

H+ + OH + H2O

  • Then:

1\, mol\, NaOH=1\, mol\, OH^{-}\Rightarrow 1\, mol\, H^{+}

1.38x10^{-3}\, mol\, OH^{-}\Rightarrow \mathbf{1.38x10^{-3}\, mol\, H^{+}}

  • In addition, one molecule of unknown acid has 2 reactive protons, then

2 mol H+ = 1 mol unknown acid

1.38x10^{-3}\, mol\, H^{+} \Rightarrow \mathbf{6.9x10^{-4}\, mol\, unknown\, acid}

c) Molecular weight of your unknown acid?

  • The molecular weight of unknown acid is defined as

Molecular\, Weight=\frac{mass}{moles}

  • The mass of unknown acid used is

mass\, unknown\, acid=(Weighing\, paper\, \&\, unknown\, acid)-(Weighing\, paper)

mass\, unknown\, acid=(0.4751\, g)-(0.3958\, g)=\mathbf{0.0793\, g}

  • Then,

Molecular\, Weight=\frac{0.0793\, g}{6.9x10^{-4}\, mol}=\mathbf{114.93\, g/mol}​​​​​​​